# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ Abstract
Quantifying physical concepts in terms of the ultimate performance of a given task has been central to theoretical progress, as illustrated by thermodynamic entropy and entanglement entropy, which respectively quantify irreversibility and quantum correlations. Symmetry breaking is equally universal, yet lacks such an operational quantification. While an operational characterization of symmetry breaking through asymptotic state-conversion efficiency is a central goal of the resource theory of asymmetry (RTA), such a characterization has so far been completed only for the $U(1)$ group among continuous symmetries. Here, we identify the complete measure of symmetry breaking for a general continuous symmetry described by any compact Lie group. Specifically, we show that the asymptotic conversion rate between many copies of pure states in RTA is determined by the quantum geometric tensor, thereby establishing it as the complete measure of symmetry breaking. As an immediate consequence of our conversion rate formula, we also resolve the Marvian-Spekkens conjecture on conditions for reversible conversion in RTA, which has remained unproven for over a decade. By applying our analysis to a standard setup in quantum thermodynamics, we show that asymptotic state conversion under thermal operations generally requires macroscopic coherence in the thermodynamic limit.
@@ I Introduction
Quantifying fundamental concepts through the ultimate performance of a given task has driven theoretical advances in physics. Thermodynamic entropy, emerging from the study of the ultimate limits of state transformation under adiabatic operations, serves as a measure of irreversibility and has led to a modern formulation of the second law of thermodynamics [1]. Likewise, entanglement entropy was discovered as a measure of quantum correlations that fully characterizes the constraints on state conversion imposed by local operations and classical communication (LOCC) [2]. Its relevance now extends beyond quantum information science, playing a pivotal role in developments across various fields of physics, including condensed matter physics [3, 4] and high-energy physics [5].
However, symmetry and its breaking, equally universal to irreversibility and quantum correlations, still lack such an operational characterization. They are cornerstones of physics, which play a vital role in characterizing natural phenomena across almost every modern field. Considering its ubiquity, quantifying symmetry breaking in the presence of symmetry in terms of operational limits will be a foundation to yield profound theoretical insights and a far-reaching impact, paralleling the influential role of entanglement entropy in modern physics.
@@
The resource theory of asymmetry (RTA) [11, 12, 13, 14, 15, 16], previously referred to as the resource theory of quantum reference frames [11, 12, 13], provides a rigorous framework for quantifying symmetry breaking. However, despite active research in RTA for over a decade [11, 12, 13, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 28, 30, 34, 35, 15, 36, 30, 14, 16, 37], progress toward identifying a complete measure for symmetry breaking remains limited. In particular, for continuous symmetries, a complete measure has been established only for the simplest $U(1)$ group [12, 23, 24], while nothing is currently known about any other continuous groups beyond a few highly specific examples of states [12, 38]. Consequently, while the RTA for the $U(1)$ group has led to applications in a wide range of fields, including quantum thermodynamics [20, 21, 23, 22], measurements [34, 35, 15, 36, 30, 39], quantum computing [32, 33, 28, 30], error-correcting codes [25, 26, 27, 28, 29, 30, 31], and black hole physics [28, 30], it has a limited impact on research domains dictated by non-Abelian Lie groups, such as $SU(2)$ spin‑rotation symmetry in quantum magnets [40, 41] and cold‑atom Hubbard models [42], and the $SU(4)$ symmetry realized in multicomponent fractional quantum Hall systems in graphene [43]. Therefore, identifying fundamental measures for symmetry breaking for non-Abelian Lie groups is an urgent task to open new avenues in these areas.
@@
Our result also provides a fundamental formula for the exploitation of symmetry breaking as a resource in quantum technologies. Concretely, it presents the ultimate limits of distillation [50], i.e., converting many low-quality resource states into a few high-quality ones, which has played a critical role in the field of entanglement, as illustrated in applications such as quantum cryptography [51], distributed quantum computing [52], and quantum internet [53]. Moreover, since our formula is expressed through a finite set of matrix inequalities, it yields a computable prediction of the fundamental efficiency bound of the distillation processes. As an immediate consequence of the formula, we also derive a necessary and sufficient condition for reversible conversion, which resolves a long-standing open problem known as the Marvian-Spekkens conjecture [14, 16].
The development of tools for quantifying symmetry breaking in RTA enables applications across diverse scenarios where dynamics are constrained by conservation laws. A particularly important case is energy conservation, which is equivalently expressed as time-translation symmetry. The breaking of this symmetry—namely, coherent superpositions of energy eigenstates with different eigenvalues—represents a fundamental resource, commonly referred to as energy coherence. Quantum thermodynamics [10, 20, 54, 55, 56, 57, 58] offers a natural setting for studying this resource, as its central objective is to determine which state transformations are achievable under thermal operations—evolutions driven by interactions with a heat bath that preserve total energy. For states without any energy coherence, called quasiclassical states, the non-equilibrium free energy fully characterizes state convertibility [10]. However, the situation for general, non-quasiclassical states remains unsettled [57, 58]. The difficulty arises from the fact that thermal operations cannot generate energy coherence, which must therefore be regarded as an independent resource, distinct from non-equilibrium free energy.
Existing approaches [10, 57, 58] have often quantified the required coherence in terms of the energy range of an external system that provides energy coherence, concluding that only a small source of coherence is needed in the thermodynamic limit. However, this perspective does not provide a refined characterization of state-specific resource requirements, as it reflects only the specifications of the external system. Building on the establishment of the QGT as a complete measure of asymmetry for a broad class of continuous symmetries, we revisit this problem. We demonstrate that, contrary to the conventional view, certain state conversions via thermal operations require a macroscopic amount of asymmetry—including energy coherence—that scales extensively with system size when quantified with the QGT. This finding highlights that the choice of asymmetry measure is a decisive factor for advancing our understanding of resource requirements in thermodynamic processes.
The significance of quantifying symmetry breaking is not limited to the study of the dynamics of symmetric systems, but it also extends to the investigation of quantum reference frames. Any physical operation is fundamentally defined relative to a reference structure, such as a time origin or a Cartesian frame. If such a structure is absent or misaligned, operations cannot, in general, be perfectly implemented. This effective lack of reference gives rise to superselection rules [11], which restrict the set of feasible operations. A symmetry-breaking quantum state therefore serves as a valuable resource for overcoming these restrictions. For example, a state that is not invariant under time translation can serve as a quantum clock, specifying the time origin. Similarly, a state that is not invariant under spatial rotations can function as a quantum gyroscope, defining the orientation of a Cartesian frame. Building on this perspective, the RTA has in fact been developed as a theory of quantum reference frames in earlier works [11, 12, 13].
@@
The rest of this paper is organized as follows: We first review the framework of RTA and the definition of the QGT in Sec. II. The main result of this paper, i.e., the approximate asymptotic conversion rate among pure states in RTA for an arbitrary Lie group expressed in terms of QGTs, is presented in Sec. III. We also apply our formula for RTA for several setups, providing a unified understanding of prior studies on conversion rates in RTA for different groups [12, 14, 16, 38, 24, 59], and proves the Marvian-Spekkens conjecture [16, 14]. In Sec. IV, we show that the QGT is a resource measure in the RTA for pure states under one-shot and asymptotic conversion. As a consequence, we prove an inequality providing a bound for the conversion rate. In Sec. V, we show the optimality of this bound, which completes the proof of the main result. In Sec. VI, we show that a sublinear number of additional resource states can significantly improve the conversion rate. In Sec. VII, we extend our analysis to mixed states to obtain the limitations of dilution and distillation processes. In Sec. VIII, as an application of our results, we investigate resource requirements in quantum thermodynamics with multiple conserved quantities, showing that certain state conversions require a macroscopic amount of asymmetry, including coherence, in the thermodynamic limit. In Sec. IX, we review the connection between symmetry breaking and quantum reference frames, and show that our main result enables us to introduce a standardized reference state of reference frames.
@@ II.1 Resource theory of asymmetry
Just like in other resource theories, including the entanglement theory, RTA is defined by specifying the free states and free operations that can be freely prepared and implemented. In RTA, free states are symmetric states, which are invariant under any symmetry transformations. That is, a state $\rho$ is $G$-symmetric iff $U(g)\rho U(g)^{\dagger}=\rho$ for all $g\in G$. Free operations are $G$-covariant channels. Here, a channel $\mathcal{E}$ is $G$-covariant iff it satisfies $\mathcal{E}\circ\mathcal{U}_{g}=\mathcal{U}_{g}^{\prime}\circ\mathcal{E}$ for all $g\in G$, where $\mathcal{U}_{g}$ and $\mathcal{U}_{g}^{\prime}$ are defined by $\mathcal{U}_{g}(\cdot)\coloneqq U(g)(\cdot)U(g)^{\dagger}$ and $\mathcal{U}_{g}^{\prime}(\cdot)\coloneqq U^{\prime}(g)(\cdot)U^{\prime}(g)^{\dagger}$ using projective unitary representations $U,U^{\prime}$ of $G$ on the input and output systems. Throughout this paper, we use the prime symbol to represent a quantity related to the output system.
@@
$\mathrm{Sym}_{G}(\rho)\coloneqq\{g\in G\mid U(g)\rho U(g)^{\dagger}=\rho\}$ (1)
for the input state $\rho$, and similarly for the output state $\sigma$. It is shown in [14] that if a state $\rho$ can be convertible to a state $\sigma$ via some $G$-covariant operation without error, the symmetry subgroups must satisfy $\mathrm{Sym}_{G}(\rho)\subset\mathrm{Sym}_{G}(\sigma)$.
We analyze the i.i.d. setup in RTA where identical copies of a state are converted to identical copies of another state. In this setup, we use the tensor product of a representation, which corresponds to the conservation of global quantities such as total particle number in the $U(1)$ case and total spins in the $SU(2)$ case. Concretely, we say that a channel $\mathcal{E}$ from $N$ copies of input state to $M$ copies of output state is $G$-covariant if $\mathcal{E}$ satisfies $\mathcal{E}\circ\mathcal{U}_{g}^{\otimes N}=\mathcal{U}_{g}^{\prime\otimes M}\circ\mathcal{E}$ for all $g\in G$, where $\mathcal{U}_{g}(\cdot)\coloneqq U(g)(\cdot)U(g)^{\dagger}$ and $\mathcal{U}_{g}^{\prime}(\cdot)\coloneqq U^{\prime}(g)(\cdot)U^{\prime}(g)^{\dagger}$.
We finally define the asymptotic conversion rate. We say that a state $\rho$ is approximately asymptotically convertible to another state $\sigma$ with a conversion rate $r$ iff there exists a sequence $\{\mathcal{E}_{N}\}_{N}$ of $G$-covariant channels such that $\lim_{N\to\infty}T\left(\mathcal{E}_{N}(\rho^{\otimes N}),\sigma^{\otimes\lfloor rN\rfloor}\right)=0$ with the trance distance $T$, which has operational significance in state distinguishability [60, 61]. We denote this conversion by
@@
and we define the approximate asymptotic conversion rate $R(\rho\to\sigma)$ by the supremum of achievable conversion rates $r$. In the following, we simply call $R(\rho\to\sigma)$ conversion rate. In this paper, we establish a formula for calculating this conversion rate among pure states, which is applicable to any compact Lie group $G$.

### v4

# Quantum geometric tensor determines the pure-state i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

@@ Abstract
Quantifying physical concepts in terms of the ultimate performance of a given task has been central to theoretical progress, as illustrated by thermodynamic entropy and entanglement entropy, which respectively quantify irreversibility and quantum correlations. Symmetry breaking is equally universal, yet lacks such an operational quantification. While an operational characterization of symmetry breaking through asymptotic state-conversion efficiency is a central goal of the resource theory of asymmetry (RTA), such a characterization has so far been completed only for the $U(1)$ group among continuous symmetries. Here, we identify the complete measure of symmetry breaking for a general continuous symmetry described by any compact Lie group. Specifically, we show that the asymptotic conversion rate between many copies of pure states in RTA is determined by the quantum geometric tensor, thereby establishing it as the complete measure of symmetry breaking. As an immediate consequence of our conversion rate formula, we also resolve the Marvian-Spekkens conjecture on conditions for reversible conversion in RTA, which has remained unproven for over a decade. Leveraging the connection between symmetry breaking and the theory of quantum reference frames, we also systematically introduce a standardized reference state for frameness based on our asymptotic conversion theory. In addition, by applying our analysis to a standard quantum-thermodynamic scenario, we show that asymptotic state conversion in contact with heat baths generally requires macroscopic coherence in the thermodynamic limit.
@@ I Introduction
Quantifying fundamental concepts through the ultimate performance of a given task has driven theoretical advances in physics. Thermodynamic entropy, emerging from the study of the ultimate limits of state transformation under adiabatic operations, serves as a measure of irreversibility and has led to a modern formulation of the second law of thermodynamics [1]. Likewise, entanglement entropy was discovered as a measure of quantum correlations that fully characterizes the state convertibility under local operations and classical communication (LOCC) [2]. Its relevance now extends beyond quantum information science, playing a pivotal role in developments across various fields of physics, including condensed matter physics [3, 4] and high-energy physics [5].
However, symmetry and its breaking, equally universal to irreversibility and quantum correlations, still lack such an operational characterization. They are cornerstones of physics, which play a vital role in characterizing natural phenomena across almost every modern field. Considering its ubiquity, quantifying symmetry breaking under symmetry constraints in terms of operational limits will be a foundation to yield profound theoretical insights and a far-reaching impact, paralleling the influential role of entanglement entropy in modern physics.
@@
The resource theory of asymmetry (RTA) [11, 12, 13, 14, 15, 16], previously referred to as the resource theory of quantum reference frames [11, 12, 13], provides a rigorous framework for quantifying symmetry breaking. However, despite active research in RTA for over a decade [11, 12, 13, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 30, 34, 35, 15, 36, 30, 14, 16, 37], progress toward identifying a complete measure for symmetry breaking remains limited. In particular, for continuous symmetries, a complete measure has been established only for the simplest $U(1)$ group [12, 23, 24], while nothing is currently known about any other continuous groups beyond a few highly specific examples of states [12, 38]. Consequently, while the RTA for the $U(1)$ group has led to applications in a wide range of fields, including quantum thermodynamics [20, 21, 23, 22], measurements [34, 35, 15, 36, 30, 39], quantum computing [32, 33, 28, 30], error-correcting codes [25, 26, 27, 28, 29, 30, 31], and black hole physics [28, 30], it has a limited impact on research domains dictated by non-Abelian Lie groups, such as $SU(2)$ spin‑rotation symmetry in quantum magnets [40, 41] and cold‑atom Hubbard models [42], and the $SU(4)$ symmetry realized in multicomponent fractional quantum Hall systems in graphene [43]. Therefore, identifying fundamental measures for symmetry breaking for non-Abelian Lie groups is an urgent task to open new avenues in these areas.
@@
Our result also provides a fundamental formula for the exploitation of symmetry breaking as a resource in quantum technologies. Concretely, it presents the ultimate limits of distillation [50], i.e., converting many low-quality resource states into a few high-quality ones, which has played a critical role in the field of entanglement, as illustrated in applications such as quantum cryptography [51], distributed quantum computing [52], and quantum internet [53]. Moreover, since our formula is expressed through a single matrix inequality, it yields a computable prediction of the fundamental efficiency bound of the distillation processes. As an immediate consequence of the formula, we also derive a necessary and sufficient condition for reversible conversion, which resolves a long-standing open problem known as the Marvian-Spekkens conjecture [14, 16].
The significance of quantifying symmetry breaking also extends to the investigation of quantum reference frames. Any physical operation is fundamentally defined relative to a reference structure, such as a time origin or a Cartesian frame. If such a structure is absent or misaligned, operations cannot, in general, be perfectly implemented. This effective lack of reference gives rise to superselection rules [11], which restrict the set of feasible operations. A symmetry-breaking quantum state therefore serves as a valuable resource for overcoming these restrictions. For example, a state that is not invariant under time translation can serve as a quantum clock, specifying the time origin. Similarly, a state that is not invariant under spatial rotations can function as a quantum gyroscope, defining the orientation of a Cartesian frame. Building on this perspective, the RTA has in fact been developed as a theory of quantum reference frames in earlier works [11, 12, 13].
@@
The development of tools for quantifying symmetry breaking in RTA enables applications across diverse scenarios where dynamics are constrained by conservation laws. A particularly important case is energy conservation, which is equivalently expressed as time-translation symmetry. The breaking of this symmetry—namely, coherent superpositions of energy eigenstates with different eigenvalues—represents a fundamental resource, commonly referred to as energy coherence. Quantum thermodynamics [10, 20, 54, 55, 56, 57, 58] offers a natural setting for studying this resource, as its central objective is to determine which state transformations are achievable in a system in contact with thermal baths under total energy conservation.
In the restricted setting where only a single heat bath at a fixed temperature is available, the associated transformations are termed thermal operations, which are central to the resource theory of athermality [10, 20, 54, 55, 56, 57, 58] (see also [59, 60, 61, 22, 62, 63] for recent works in this area). For states without any energy coherence, called quasiclassical states, the non-equilibrium free energy fully characterizes state convertibility under thermal operations [10]. However, the situation for general, non-quasiclassical states remains unsettled [57, 58]. The difficulty arises from the fact that thermal operations cannot generate energy coherence, which must therefore be regarded as an independent resource, distinct from non-equilibrium free energy. Existing approaches [10, 57, 58] have often quantified the required coherence in terms of the operator norm of Hamiltonian of an external system that provides energy coherence, suggesting that only a small source of coherence is needed in the thermodynamic limit. However, this perspective does not yield a refined characterization of state-specific resource requirements, as it reflects only the external system’s specifications.
Building on the establishment of the QGT as a complete measure of asymmetry for a broad class of continuous symmetries, we revisit this problem in a more general scenario involving contact with multiple baths at different temperatures under energy conservation. We demonstrate that, contrary to the conventional view, achieving certain state conversions via thermal contact require a macroscopic amount of energy coherence that scales extensively with system size when quantified with the QGT, even when baths at different temperatures are allowed. This reveals that macroscopic coherence can be necessary in quantum thermodynamics not only in the single-bath setting, but also in multi-bath scenarios relevant, for example, to heat-engine setups. This finding highlights that the choice of asymmetry measure is a decisive factor for advancing our understanding of resource requirements in thermodynamic processes. Moreover, we show that the same observation extends to settings with additional conserved charges, such as particle number or angular momentum, implying that state transformations in thermodynamic processes generally require an extensive amount of asymmetry.
Outline
The outline of the remainder of the paper with a summary of the main results and key ideas are as follows.
In the Preliminaries (Sec. II), we review and introduce key notions required to state our main results. Specifically,
  • In Sec. II.1, we review the standard framework of the RTA. We also introduce the conversion rate between states in the RTA, which is the primary quantity we study in this work.
  • In Sec. II.2, we briefly review the QGT, which quantifies the sensitivity of a quantum state under an infinitesimal change of parameters. In our study of symmetries associated with a compact Lie group $G$, the QGT is a $\dim G\times\dim G$ positive semidefinite matrix, which can be expressed as a non-symmetrized covariance matrix of the generators of infinitesimal group transformations (Eq. (7)).
In Sec. III, we first present the main theorem of this paper (Theorem 1). Equation (8) gives an explicit formula for the conversion rate between i.i.d. pure states, expressed as a single matrix inequality between the QGTs of the input and output states. This establishes the QGT as the complete measure of asymmetry for continuous symmetries described by a compact Lie group. We further rewrite the formula in terms of the quantum max-relative entropy between the QGTs, which makes its computability via semidefinite programming transparent. Finally, we illustrate the formula for various symmetry groups and clarify its relation to prior studies. Specifically,
  • In Sec. III.1, we apply the formula to the $U(1)$ symmetry, which reproduces the result of [24].
  • In Sec. III.2, as a corollary of the main result, we prove a necessary and sufficient condition for asymptotically reversible conversion for a compact Lie group (Corollary 2). We then show that this result proves the Marvian–Spekkens conjecture [16, 14], which has remained open for over a decade. This also yields reversible conversion rates for specific examples that were calculated in earlier, separate studies [12, 38, 24].
  • In Sec. III.3, we study the irreversibility of asymptotic conversion. As an explicit example, we compute the conversion rate for the $SU(2)$ group, obtaining a result consistent with a prior study for a restricted class of pure states [12].
  • In Sec. III.4, we show that our formula also applies to finite groups. Consequently, we find that the conversion rate is either infinite or zero, consistent with [64].
In Sec. IV, we explain the key ideas used to prove the main theorem. The proof consists of two parts: the converse part, which establishes an upper bound on the conversion rate, and the direct part, which proves the achievability of this bound (i.e., optimality). Specifically,
  • In the converse part (Sec. IV.1), we first prove the monotonicity of the QGT. To this end, we first establish a connection between the QGT and Petz’s monotone metrics [65, 66]. Using the monotonicity of Petz’s metrics, we show that the QGT is non-increasing under exact conversion, in the sense of a matrix inequality, i.e., the Löwner order. This implies that the QGT is a valid asymmetry monotone (Eq. (31)).
    Since the QGT is additive, it is natural to expect that its linear rate (i.e., the QGT per copy) provides a useful monotone in the asymptotic regime (analogous to densities of extensive quantities in the thermodynamic limit). Proving monotonicity of this QGT rate for asymptotic conversion with vanishing error, however, is technically delicate. The key difficulty is that the QGT can, in general, scale quadratically with system size, so its linear rate can change drastically even under small conversion errors. This phenomenon is known as asymptotic discontinuity [67, 68, 13, 23, 24, 69]. Nonetheless, by carefully analyzing the asymptotic behavior of Petz’s monotone metrics, we establish monotonicity of the QGT rate (Eq. (37)). This monotonicity in turn yields an upper bound on the conversion rate (Eq. (38)). The technical details of the converse part are given in Appendix C.
  • In the direct part (Sec. IV.2), we prove the achievability of the above bound (and hence optimality). To this end, we explicitly construct conversion channels. A key observation is that convertibility between states in the RTA is directly related to convertibility between quantum statistical models (Lemma 4). The asymptotic behavior of these statistical models is analyzed using quantum local asymptotic normality [70, 71, 72, 73], which identifies a limit model that approximates the behavior of i.i.d. statistical models in the asymptotic regime. Importantly, we find that this limit model is characterized by the QGT associated with the underlying statistical model. Building on this observation, we construct the conversion channels via what we call an “estimate-and-convert” strategy. The technical details of the direct part are given in Appendix D.
Having established the pure-state theory of asymptotic conversion, we next extend our analysis to mixed states in Sec. V. We consider two fundamental scenarios: asymmetry distillation and asymmetry dilution. Although these results do not provide a complete characterization of mixed-state conversion rates, they serve as a foundation for future work. Specifically,
  • In Sec. V.1, we study the conversion rate from i.i.d. mixed states to i.i.d. pure states in the RTA, namely, the distillable asymmetry. To this end, we introduce an extension of the QGT to mixed states and show that it is a valid asymmetry monotone. Consequently, we derive an upper bound on the distillable asymmetry and obtain a simple condition under which the distillable asymmetry vanishes (Corollary 6). When specialized to time-translation symmetry, this condition reduces to the result of [23].
  • In Sec. V.2, we study the opposite scenario, namely, the conversion rate from i.i.d. pure states to i.i.d. mixed states in the RTA, i.e., the asymmetry cost. By combining the standard typical-sequence argument [74, 24] with our estimate-and-convert strategy, we prove an upper bound on the asymmetry cost by explicitly constructing conversion channels (Proposition 7). This bound is tight for $U(1)$ symmetry, reproducing the result of [24]. While we leave open whether the bound is tight for a general Lie group, we clarify why the $U(1)$ argument does not directly extend to the general case.
Building on the results above, in Sec. VI we apply our analysis to quantum reference frames. We begin with a careful review of the relation between asymmetry and quantum reference frames. In general, any physical operation is defined relative to some reference structure, such as a clock or a spatial direction. In the absence of an absolute (i.e., shared) reference frame, asymmetry becomes the resource required to implement such operations. In this sense, asymmetry can also be regarded as a resource of frameness. Specifically,
  • We show that our pure-state conversion theory enables us to introduce a standardized reference state that serves as a benchmark for comparing frameness, analogous to the role of an ebit in entanglement theory. So far, such a standardized reference state has been proposed only for the $SU(2)$ group [38]. Our conversion theory provides a systematic way to define standardized reference states for any semisimple compact Lie group. Concretely, we define a standardized reference state in Eq. (61). This state is a universal frameness resource, in the sense that it can be converted into any pure state at a nonzero rate. As a consequence of the conversion-rate formula, we further show that the rate is given by the min-entropy of the QGT (Eq. (65)).
  • We then clarify the distinction between differential-geometric quantities (such as the QGT and the quantum Fisher information) and commonly used entropic quantities (such as $G$-asymmetry) from the viewpoint of quantum reference frames. For instance, for the rotation group, the former quantifies both the direction and the magnitude of asymmetry, whereas the latter is insensitive to direction. Therefore, if we are concerned not only with the amount of asymmetry but also with the direction in which a state breaks the symmetry, geometric quantities provide more refined information.
In addition, we apply our framework to thermodynamics by considering interactions with thermal baths under conservation laws. The central physical setting is described in Sec. VII.1: a system couples to heat baths initially prepared in Gibbs states via a global unitary that conserves the total energy. When all baths are at the same fixed temperature, the resulting dynamics reduces to thermal operations. Here, however, we allow the baths to be at different temperatures, so that our framework applies to a broader class of thermodynamic scenarios. We further extend this setting to include additional conserved quantities, such as particle number and angular momentum, associated with a Lie group.
Using the tools developed in this work, we aim to demonstrate nontrivial constraints on state transformations under such thermal contact and to analyze the resources required to circumvent them. To this end, we relate thermal contact to covariant operations in Sec. VII.2. A key theoretical ingredient is Theorem 8, which establishes monotonicity of the QGT in the asymptotic distillation of i.i.d. pure states. Importantly, this theorem applies to general initial states of the system (including non-i.i.d. and mixed states), enabling analysis of a wide range of scenarios.
In particular, we study the following two classes of scenarios:
  • In Sec. VII.3, we consider the asymptotic distillation of i.i.d. pure states from general i.i.d. states. Using the monotonicity of the QGT, we establish a no-go theorem for distillation (Corollary 9). In particular, we show that for a typical full-rank state, no asymmetric pure state can be distilled under thermal contact at a linear rate. We illustrate this result with the example of $U(1)$ symmetry.
  • In Sec. VII.4, we consider a generalized asymptotic distillation setting that allows for an external resource system assisting the conversion. By applying the monotonicity of the QGT (Theorem 8), we show that circumventing the above no-go theorem requires the external resource system to possess a macroscopic amount of asymmetry. In the case of time-translation symmetry, this implies that asymptotic state transformation under thermal contact in general requires macroscopic energy coherence (Eq. (96)).
Finally, we present our conclusions in Sec. VIII.
@@ II.1 Resource theory of asymmetry
Just like in other resource theories, including the entanglement theory, RTA is defined by specifying the free states and free operations that are considered to be freely prepared and implemented. In RTA, free states are symmetric states, which are invariant under any symmetry transformations. That is, a state $\rho$ is $G$-symmetric iff $U(g)\rho U(g)^{\dagger}=\rho$ for all $g\in G$. Free operations are $G$-covariant channels. Here, a channel $\mathcal{E}$ is $G$-covariant iff it satisfies $\mathcal{E}\circ\mathcal{U}_{g}=\mathcal{U}_{g}^{\prime}\circ\mathcal{E}$ for all $g\in G$, where $\mathcal{U}_{g}$ and $\mathcal{U}_{g}^{\prime}$ are defined by $\mathcal{U}_{g}(\cdot)\coloneqq U(g)(\cdot)U(g)^{\dagger}$ and $\mathcal{U}_{g}^{\prime}(\cdot)\coloneqq U^{\prime}(g)(\cdot)U^{\prime}(g)^{\dagger}$ using projective unitary representations $U$ and $U^{\prime}$ of $G$ on the input and output systems, respectively. Throughout this paper, we use the prime symbol to represent a quantity related to the output system.
The physical implication of $G$-covariance is clarified by the covariant Stinespring dilation theorem [75, 14], which states that a $G$-covariant channel admits a representation as the dynamics of an open system interacting with an environment under conservation laws. More precisely, a quantum channel $\mathcal{E}$ is $G$-covariant if and only if there exist ancillary systems $E$ and $E^{\prime}$, carrying projective unitary representations $U_{E}$ and $U_{E^{\prime}}$ of $G$, such that
$\mathcal{E}(\cdot)=\mathrm{Tr}_{E^{\prime}}\left(V\left(\cdot\otimes\eta_{E}\right)V^{\dagger}\right),$ (1)
where $\eta_{E}$ is a $G$-symmetric state on $E$, and $V$ is a unitary operator satisfying
$V(U(g)\otimes U_{E}(g))=(U^{\prime}(g)\otimes U_{E^{\prime}}(g))V,\quad\forall g\in G.$ (2)
Since Eq. (2) implies that the joint unitary evolution $V$ respects the symmetry described by $G$, a $G$-covariant channel characterizes the most general open-system dynamics consistent with conservation laws, without supplying asymmetric states from the environment (see also Appendix A).
@@
$\mathrm{Sym}_{G}(\rho)\coloneqq\{g\in G\mid\mathcal{U}_{g}(\rho)=\rho\}$ (3)
for the input state $\rho$, and similarly for the output state $\sigma$. It is shown in [14] that if a state $\rho$ can be converted to a state $\sigma$ via some $G$-covariant operation without error, the symmetry subgroups must satisfy $\mathrm{Sym}_{G}(\rho)\subset\mathrm{Sym}_{G}(\sigma)$.
We analyze the i.i.d. setup in RTA where identical copies of a state are converted to identical copies of another state. In this setup, we use the tensor product of a representation, which corresponds to the conservation of global quantities such as total particle number in the $U(1)$ case and total spins in the $SU(2)$ case. Concretely, we say that a channel $\mathcal{E}$ from $N$ copies of input state to $M$ copies of output state is $G$-covariant if $\mathcal{E}$ satisfies $\mathcal{E}\circ\mathcal{U}_{g}^{\otimes N}=\mathcal{U}_{g}^{\prime\otimes M}\circ\mathcal{E}$ for all $g\in G$.
We finally define the asymptotic conversion rate. We say that a state $\rho$ is asymptotically convertible to another state $\sigma$ with a conversion rate $r$ iff there exists a sequence $\{\mathcal{E}_{N}\}_{N}$ of $G$-covariant channels such that $\lim_{N\to\infty}T\left(\mathcal{E}_{N}(\rho^{\otimes N}),\sigma^{\otimes\lfloor rN\rfloor}\right)=0$. Here $T(\rho,\sigma)\coloneqq\frac{1}{2}\|\rho-\sigma\|_{1}$ is the trace distance, which has operational significance in state distinguishability [76, 77]. We denote this conversion by
@@
and we define the asymptotic conversion rate $R(\rho\to\sigma)$ by the supremum of achievable conversion rates $r$. In the following, we simply call $R(\rho\to\sigma)$ the conversion rate. In this paper, we establish a formula for calculating this conversion rate among pure states, which is applicable to any compact Lie group $G$.